# Graphusion: A RAG Framework for Scientific Knowledge Graph Construction with a Global Perspective

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: THE WEB CONFERENC; April 28–May 02, 2025; SYDNEY, AUSTRALIAisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Language modelsccs: Information systems Extraction, transformation and loadingccs: Information systems Database utilities and tools
@@
Recognizing these limitations, we propose a cost-efficient KGC approach to incorporate the global perspective and improve RAG performance on downstream question-answering (QA) tasks. Fig. 1 illustrates the differences between zero-shot LLMs and Local Graph-RAG, GraphRAG and our model. While Local Graph-RAG primarily focuses on knowledge extraction from a disconnected point of view, GraphRAG is able to summarize the information from a global view. However, our approach, Graphusion, incorporates a knowledge fusion step to integrate local knowledge into a global context directly. The core fusion step performs global merging and resolution across multiple local sub-graphs to form one large connected KG. Specifically, we leverage LLMs not only for extraction but also for critical knowledge integration, marking the first initiative to utilize LLMs for such a comprehensive merging process. Moreover, we generate a seed entity list to guide the entity extraction with better granularity. We demonstrate Graphusion’s capability in knowledge graph construction and link prediction. Our results show that Graphusion achieves scores of 2.92 and 2.37 out of 3 for entity extraction and relation recognition, respectively. Furthermore, we show that a simplified link prediction prompt outperforms supervised learning baselines, achieving a 3% higher F1 score. Then, to further show the power of the constructed KG, we showcase how it could be applied to the NLP domain, and most importantly, we validate it in an educational scenario with complex QA tasks.  [^1] https://anonymous.4open.science/r/CGPrompt-C9C7/readme.md.
@@
Problem Definition
A $KG$ is defined as a set of triplets $KG=\{(h_{i},r_{i},t_{i})\mid h_{i},t_{i}\in E,\ r_{i}\in R,\ i=1,2,\ldots,n\}$, where $E$ is the set of entities, $R$ is the set of possible relations, and $n$ is the total number of triplets in the $KG$. The task of zero-shot KGC involves taking a set of free text $T$ and generating a list of triplets $(h,r,t)$ spanning a KG. Optionally, there is an expert-annotated KG, $G_{E}$, as input, in order to provide existing knowledge. In our setting, the number of triplets of $KG$ is much larger than $G_{E}$. We select the domain to be NLP, so the entities are limited to NLP entities, with other entity types such as people, and organizations not being our focus. Referring to previous works (Luan et al., 2018), we define 7 relations types: Prerequisite_of, Used_for, Compare, Conjunction, Hyponym_of, Evaluate_for and Part_of. We will now describe the three steps of the pipeline in detail.
@@ 4. Once step 3 is done, consider every
1. Entity Quality
Excellent (3 points): Both entities are highly relevant and specific to the domain. At an appropriate level of detail, neither too broad nor too specific. For example, an entity could be introduced by a lecture slide page, or a whole lecture, or possibly have a Wikipedia page. Acceptable (2 points): Entity is somewhat relevant, or granularity is acceptable. Poor (1 point): Entity is at an inappropriate level of detail, too broad or too specific.
2. Relation Quality
Correct (3 points): The relation logically and accurately describes the relation between the head and tail entities. Somewhat Correct (2 points): The relation is acceptable but has minor inaccuracies or there might be another better or correct answer. Incorrect (1 point): The relation does not logically describe the relation between the entities.
@@ 2. Relation Quality
Case Study: Entity Extraction To demonstrate the importance of the seed entity list from Step 1, we examine a selection of random entities extracted by both GraphRAG and Graphusion, as shown in Tab. 2. All results are based on the GPT-4o backbone. GraphRAG sometimes extracts overly general terms, such as benchmark and methodology, which occur frequently in the corpus. In subsequent experiments, we will further illustrate how GraphRAG tends to extract entities with unsuitable granularity.
GraphRAG Graphusion
mixture-of-experts technique code-switching tasks
mbart NAS-BERT
benchmark linear indexed grammars
multilingual alignment analyzing data
diffusionbert semantic parsing
proposed method bias-variance
methodology few-shot learning
@@
Experimental Results Our analysis in Tab. 6 compares the results based on Graphusion and two baselines, including GPT-4o zero-shot (zs) and GPT-4o with RAG (RAG). Our method with the Graphusion constructed KG shows significant improvements across Tasks 1 to 6 over the baselines. Specifically, Task 6 is evaluated by two NLP experts, with a Kappa score of 0.67, which suggests substantial agreement. The results indicate that our pipeline exhibits a marginally superior performance, particularly in the expert evaluation of Convincity and Factuality. This suggests that our method might be better at generating content that is not only factually accurate but also presents it in a more persuasive way to the reader. Compared to the base RAG framework, the Graphusion-generated KGs lead to better performance, particularly in Task 4 and Task 5, where a global understanding is essential. This improvement highlights the critical role of our core fusion step in addressing complex QA. We include details on comparing with GraphRAG in the Appendix.
Case Study: Task 2 & Task 3 (Entity counts) As depicted in Fig  5, we evaluate the average number of entities created by GPT-4o zs and our Graphusion framework in the responses for Task 2 and Task 3, in which both tasks require the model to give a list of reasonable entities. The results show that without the enhancement of KG retrieved information, GPT-4o tends to mention more entities in the generated response (Task 2: 11.04, Task 3: 11.54), which might be irrelevant or broad.
Figure 5. Entity counts in Task 2 and Task 3.
@@ Conclusion
Ethical Statement
We encourage human verification to ensure accuracy in automated knowledge extraction, especially in critical fields such as education. Unlike pure LLM applications and LLM RAG applications, the knowledge stored in KGs can be manually verified, which can enhance the trustworthiness of AI applications that use KGs to retrieve background knowledge. To address biases in LLMs and KGs, we used trusted data sources, but LLM biases may persist. No personal or sensitive data were used, and all experiments followed ethical AI standards.
@@ 1. Identify all entities: For each identified entity, extract the following information:
Appendix B Experimental Setup
@@ Appendix C TutorQA
Appendix D Knowledge Graph Construction Analysis

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: TBDconference: ; April 28-May 2, 2025; Sydney, NSW, Australiaisbn: TBDccs: Information systems Language modelsccs: Information systems Extraction, transformation and loadingccs: Information systems Database utilities and tools
Introduction
@@
Recognizing these limitations, we propose a cost-efficient KGC approach to incorporate the global perspective and improve RAG performance on downstream question-answering (QA) tasks. Fig. 1 illustrates the differences between zero-shot LLMs and Local Graph-RAG, GraphRAG and our model. While Local Graph-RAG primarily focuses on knowledge extraction from a disconnected point of view, GraphRAG is able to summarize the information from a global view. However, our approach, Graphusion, incorporates a knowledge fusion step to integrate local knowledge into a global context directly. The core fusion step performs global merging and resolution across multiple local sub-graphs to form one large connected KG. Specifically, we leverage LLMs not only for extraction but also for critical knowledge integration, marking the first initiative to utilize LLMs for such a comprehensive merging process. Moreover, we generate a seed entity list to guide the entity extraction with better granularity. We demonstrate Graphusion’s capability in knowledge graph construction and link prediction. Our results show that Graphusion achieves scores of 2.92 and 2.37 out of 3 for entity extraction and relation recognition, respectively. Furthermore, we show that a simplified link prediction prompt outperforms supervised learning baselines, achieving a 3% higher F1 score. Then, to further show the power of the constructed KG, we showcase how it could be applied to the NLP domain, and most importantly, we validate it in an educational scenario with complex QA tasks.  [^1] https://github.com/IreneZihuiLi/Graphusion
@@
Problem Definition A $KG$ is defined as a set of triplets $KG=\{(h_{i},r_{i},t_{i})\mid h_{i},t_{i}\in E,\ r_{i}\in R,\ i=1,2,\ldots,n\}$, where $E$ is the set of entities, $R$ is the set of possible relations, and $n$ is the total number of triplets in the $KG$. The task of zero-shot KGC involves taking a set of free text $T$ and generating a list of triplets $(h,r,t)$ spanning a KG. Optionally, there is an expert-annotated KG, $G_{E}$, as input, in order to provide existing knowledge. In our setting, the number of triplets of $KG$ is much larger than $G_{E}$. We select the domain to be NLP, so the entities are limited to NLP entities, with other entity types such as people, and organizations not being our focus. Referring to previous works (Luan et al., 2018), we define 7 relations types: Prerequisite_of, Used_for, Compare, Conjunction, Hyponym_of, Evaluate_for and Part_of. We will now describe the three steps of the pipeline in detail.
@@ 4. Once step 3 is done, consider every
1. Entity Quality Excellent (3 points): Both entities are highly relevant and specific to the domain. At an appropriate level of detail, neither too broad nor too specific. For example, an entity could be introduced by a lecture slide page, or a whole lecture, or possibly have a Wikipedia page. Acceptable (2 points): Entity is somewhat relevant, or granularity is acceptable. Poor (1 point): Entity is at an inappropriate level of detail, too broad or too specific.
2. Relation Quality Correct (3 points): The relation logically and accurately describes the relation between the head and tail entities. Somewhat Correct (2 points): The relation is acceptable but has minor inaccuracies or there might be another better or correct answer. Incorrect (1 point): The relation does not logically describe the relation between the entities.
@@
Experimental Results Our analysis in Tab. 5 compares the results based on Graphusion and two baselines, including GPT-4o zero-shot (zs) and GPT-4o with RAG (RAG). Our method with the Graphusion constructed KG shows significant improvements across Tasks 1 to 6 over the baselines. Specifically, Task 6 is evaluated by two NLP experts, with a Kappa score of 0.67, which suggests substantial agreement. The results indicate that our pipeline exhibits a marginally superior performance, particularly in the expert evaluation of Convincity and Factuality. This suggests that our method might be better at generating content that is not only factually accurate but also presents it in a more persuasive way to the reader. Compared to the base RAG framework, the Graphusion-generated KGs lead to better performance, particularly in Task 4 and 5, where a global understanding is essential. This improvement highlights the critical role of our core fusion step in addressing complex QA.
@@
Extension on Japanese Medical Data
In our exploration of extending Graphusion to Japanese medical data, we utilized a dataset comprising approximately 0.1 billion tokens collected from Japanese drug instructions through data crawling (Qiu et al., 2024). Example triplets generated by Graphusion include: (バラシクロビル錠500mg, 抑制される, 発疹) ((Valacyclovir Tablets 500mg, suppressed, rash)), (ゾビラックス錠400, 作用機序,ウイルスDNAの複製を阻害することによりウイルスの増殖を抑える) ((Zovirax Tablets 400, mechanism of action, inhibits the replication of viral DNA to suppress the proliferation of the virus)). We randomly selected several case studies and found that the generated triplets were reasonable, demonstrating that Graphusion exhibits good generalizability. However, conducting a comprehensive evaluation is challenging due to the significant human effort required; therefore, we leave this evaluation for future work.
@@ 1. Identify all entities: For each identified entity, extract the following information:
Appendix B Graphusion Case Study: Entity Extraction
To demonstrate the importance of the seed entity list from Step 1, we examine a selection of random entities extracted by both GraphRAG and Graphusion, as shown in Tab. 6. All results are based on the GPT-4o backbone. GraphRAG sometimes extracts overly general terms, such as benchmark and methodology, which occur frequently in the corpus. In subsequent experiments, we will further illustrate how GraphRAG tends to extract entities with unsuitable granularity.
GraphRAG Graphusion
mixture-of-experts technique code-switching tasks
mbart NAS-BERT
benchmark linear indexed grammars
multilingual alignment analyzing data
diffusionbert semantic parsing
proposed method bias-variance
methodology few-shot learning
Appendix C Experimental Setup
@@ Appendix D TutorQA
Entity counts As depicted in Fig  6, we evaluate the average number of entities created by GPT-4o zs and our Graphusion framework in the responses for Task 2 and Task 3, in which both tasks require the model to give a list of reasonable entities. The results show that without the enhancement of KG retrieved information, GPT-4o tends to mention more entities in the generated response (Task 2: 11.04, Task 3: 11.54), which might be irrelevant or broad.
Figure 6. Entity counts in Task 2 and Task 3.
@@
Appendix E Knowledge Graph Construction Analysis